10. An analyst is trying to determine the capital structure for Big Dawg Incorporated. After some careful research, the analyst knows the following: --Big Dawg has 1.67 million shares of common stock trading today at $21.00 per share. The book value of Big Dawg common stock is $11.00 million. The cost of equity for the firm is estimated to be 11.00%. --Big Dawg has $8.00 million in long-term debt on its balance sheet. The debt is trading at 91.00% of face value in the market with a yield to maturity of 10.00%. The tax rate facing the firm is 37.00%. What portion is equity in the firm’s capital structure? Submit Answer format: Percentage Round to: 2 decimal places (Example: 9.24%, % sign required. Will accept decimal format rounded to 4 decimal places (ex: 0.0924))
Answer #1

Market value of shares = 1670000*21 = 35070000

Market value of debt = 8000000*91% = 7280000

Total capital = 42350000

Weight of equity =35070000 /42350000

= 82.81%